OpenStudy (anonymous):

Find the exact area of a circle having the given circumference 4(sqrt3pi) it is not 124!!!!

OpenStudy (anonymous):

circumference is \[2\pi r\] and here \[2\pi r = 4\sqrt{3\pi}\] so \[r=\frac{2\sqrt{3\pi}}{\pi}\] \[A=\pi r^2=\pi \times(\frac{2\sqrt{3\pi}}{\pi})^2\] \[12\pi\]

OpenStudy (anonymous):

circumference = 2r(pi) = 4(sqrt3pi) r(pi) = 2(sqrt3pi) r = 2(sqrt3pi)/pi area = pi(r^2) area = pi(2(sqrt3pi)/pi)^2 area = pi(12pi/pi^2) area = 12pi^2/pi^2 area = 12

OpenStudy (anonymous):

I'll give you a new formula to learn along with the usual two: Twice the area = radius times circumference (2A = rC).